OpenStudy (anonymous):

Metals!!!!!!! Write an inverse variation equation. A small backyard swimming pool holds 3,000 gallons of water. Derek fills the pool using a hose that delivers water at a rate of 200 gallons per hour, 225 gallons per hour, or 250 gallons per hour. For each rate x, find the number of hours y it will take to fill the pool.

OpenStudy (anonymous):

3000 / 200 = 15 hours 3000 / 225 = 13.333 hours 3000 / 250 = 12 hours???

OpenStudy (anonymous):

@Jpenrod So, the answer is 12 hours?

OpenStudy (anonymous):

yezzir